OpenStudy (anonymous):

36. The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time?

OpenStudy (anonymous):

months | inches 2 14 4 23 6 32 8 41

OpenStudy (anonymous):

"months" is your independent variable or "x" "inches" is your dependent variable or "y" First, get slope: m = (y2 - y1) / (x2 - x1) = (23 - 14) / (4 - 2) = 9/2 Next, we use the point-slope equation for a line: y - y1 = m(x - x1) -> y - 14 = (9/2)(x - 2) y - 14 = (9/2)x - 9 y = (9/2)x + 5

OpenStudy (anonymous):

That isn't one of the answers so here they are.

OpenStudy (anonymous):

y + 7 = 4/5 (x – 9) y – 7 = 4/5 (x + 9) y – 9 = 4/5 (x + 7) y + 9 = 4/5 (x – 7)

OpenStudy (anonymous):

I think you might be looking at an answer choice for a different problem there. None of those choices work and none are even in the right format. Mine is in the right format and it works for every point. Somewhere, either the problem or choice selection is not matching.

OpenStudy (anonymous):

I don't know, that is what it states.

OpenStudy (anonymous):

One more.